# Meta-Task Prompting Elicits Embedding from Large Language Models

### v1

@@ 4.2 Main Results
The results of MetaEOL on STS tasks are shown in Table 1, with a notable performance by MetaEOL which requires no training. Among models that do not require training, prompt-based methods exhibit superior results compared to average pooling methods, especially with the LLAMA and Mistral models. Across various models including LLAMA2-7B/13B and Mistral-7B, MetaEOL, which does not require any training, demonstrates competitive performance compared to contrastive-trained models such as SimCSE-BERT and SimCSE-Roberta, albeit with a slight lag behind PromptBERT. Furthermore, MetaEOL significantly outperforms PromptEOL across three test models, demonstrating a consistent improvement. Notably, the LLAMA2-13B model using MetaEOL shows an average improvement of 6.73% over PromptEOL, underscoring the efficacy of MetaEOL.
@@ 5.1 Ablation Study
We evaluate the effectiveness of key components of MetaEOL in table 3. First, to ensure the improvement observed with MetaEOL is not merely due to involving more prompts, we create seven paraphrased versions of the PromptEOL prompt, resulting in a total of eight prompts. [^2] The seven paraphrased prompts are presented in Appendix A.2 We then average the embeddings from these eight prompts to form the final sentence embedding. We find merely duplicating PromptEOL prompts (w. 8 paraphrase) does not improve PromptEOL but results in a significant decline. Additionally, we implement MetaEOL exclusively on each meta-task (TC/SA/PI/IE only). We find that tasks requiring a detailed comprehension of sentences (PI and IE) yield superior performance compared to those requiring a broader understanding, even surpassing PromptEOL. However, MetaEOL, which combines the embeddings from these meta-tasks, outperforms all individual meta-tasks, confirming the complementarity of the meta-tasks and the effectiveness of combining embeddings from diverse meta-tasks. We finally find that averaging the embeddings from different meta-tasks yields better results than concatenating them.
@@ 5.4 Influence of Output Layers
We check the impact of output layers for LLAMA2 and Mistral-7B models, using PromptEOL and MetaEOL. Figure 4 shows that the last layer is not always the most effective for STS tasks.
@@
5.5 Scaling LLMs
In this section, we investigate the impact of model size on the performance of MetaEOL. For the sake of computational resources, we load models with 4-bit precision.
The results show that using the final layer for sentence embedding generation, which is indicated by layer index -1, does not yield improved performance with increased model size. Contrastingly, the application of our proportional layer strategy reveals a different trend. Specifically, the LLAMA2-70B model, which utilizes the -8 layer, demonstrates superior performance, suggesting that larger models might benefit more significantly from selecting a proportionate layer rather than the last layer for sentence embedding. This observation could point to a potential scaling law, where larger models require a different, non-final layer to maximize performance effectively.
5.6 Transfer Learning Tasks
@@
We conclude our analysis by assessing the performance of MetaEOL on transfer learning tasks. Following prior works (Gao et al., 2021; Ni et al., 2022), we utilize the standard transfer learning tasks provided by SentEval. The tasks consist of MR (Pang and Lee, 2005), CR (Hu and Liu, 2004), SUBJ (Pang and Lee, 2004), MPQA (Wiebe et al., 2005), SST-2 (Socher et al., 2013), TREC (Voorhees and Tice, 2000), and MRPC (Dolan and Brockett, 2005). We include two supervised contrastive-trained models (SimCSE and ST5-Enc) for reference. Notably, ST5-Enc, a model with a 4.8B parameter count, is extensively trained on natural language inference (NLI) data and two billion question-answer pairs.
To investigate the ability of task-specific prompts to modify embedding behavior, we have crafted task prompts tailored to each SentEval task. [^4] The details of the task prompts are presented in Appendix A.5. As an example, for the Movie Review (MR) dataset, we designed a prompt structured as: In this task, you’re given a movie review, and you need to classify its sentiment into positive or negative. For this task, this sentence: "input sentence" means in one word:", referred to as MRPrompt in Table 6. These task-specific prompts significantly improve the corresponding task performance, always better than PromptEOL and heavily supervised contrastive-trained ST5-Enc, verifying that LLAMA2-7B can follow the prompt to generate tailored embeddings without any training. This indicates that carefully designed prompts can effectively steer the pre-trained embeddings to align with various NLP tasks, thus providing a more resource-efficient alternative to the traditional fine-tuning paradigm.
@@ 6 Conclusion
In this paper, we introduce MetaEOL, a new approach for deriving high-quality sentence embeddings from LLMs without requiring any training. By leveraging a diverse set of meta-task prompts, MetaEOL effectively captures multiple representations of sentences from distinct perspectives. We show simply averaging these meta-task derived embeddings leads to generalized general-purpose embeddings, which work remarkably well across STS datasets and transfer learning tasks. Experiments show averaging embeddings and integrating meta-tasks improves STS tasks’ performance, challenging conventional training methods.
@@
We acknowledge two limitations in our work: computational overhead and restricted evaluation benchmarks. As MetaEOL requires feeding multiple prompts to LLMs to generate several embeddings, the computational cost will be higher than previous methods. Nonetheless, in contexts where sentences are consistently reused, such as when embeddings are stored for downstream classification or retrieval tasks, this issue becomes less significant. Furthermore, our evaluation is currently confined to sentence-level tasks in English only. As LLMs continue to advance, exploring the performance of MetaEOL in multilingual contexts and its applicability to document retrieval presents an intriguing avenue for future research.
@@ A.2 Paraphrased Prompts of PromptEOL
This sentence : "input sentence" can be rephrased to one word:" This sentence : "input sentence" can be expressed as one word:" This sentence : "input sentence" implies in one word:" This sentence : "input sentence" indicates in one word:" The meaning of this sentence : "input sentence" can be conveyed in another word:" This sentence : "input sentence" can be restated as one word:" This sentence : "input sentence" can be reformulated as one word:"
@@ A.3 Prompts of MetaEOL
Text Classification General Category Identification: In this task, you’re presented with a text excerpt. Your task is to categorize the excerpt into a broad category such as ’Education’, ’Technology’, ’Health’, ’Business’, ’Environment’, ’Politics’, or ’Culture’. These categories help in organizing content for better accessibility and targeting. For this task, this sentence : "input sentence" should be classified under one general category in one word:" Opinion vs. Fact Discrimination: In this task, you’re given a statement and you need to determine whether it’s presenting an ’Opinion’ or a ’Fact’. This distinction is vital for information verification, educational purposes, and content analysis. For this task, this sentence : "input sentence" discriminates between opinion and fact in one word:" Sentiment Analysis Product Review Rating: In this task, you’re given a review from an online platform. Your task is to generate a rating for the product based on the review on a scale of 1-5, where 1 means ’extremely negative’ and 5 means ’extremely positive’. For this task, this sentence : "input sentence" reflects the sentiment in one word:" Emotion Detection: In this task, you’re reading a personal diary entry. Your task is to identify the predominant emotion expressed, such as joy, sadness, anger, fear, or love. For this task, this sentence : "input sentence" conveys the emotion in one word:" Paraphrase Identification Similarity Check: In this task, you’re presented with two sentences. Your task is to assess whether the sentences convey the same meaning. Use ’identical’, ’similar’, ’different’, or ’unrelated’ to describe the relationship. To enhance the performance of this task, this sentence : "input sentence" means in one word:" Contextual Synonym Detection: In this task, you’re given a sentence and a phrase. Your task is to determine if the phrase can be a contextual synonym within the given sentence. Options include ’yes’, ’no’, or ’partially’. To enhance the performance of this task, this sentence : "input sentence" means in one word:" Information Extraction Key Fact Identification: In this task, you’re examining a news article. Your task is to extract the most critical fact from the article. For this task, this sentence : "input sentence" encapsulates the key fact in one word:" Entity and Relation Extraction: In this task, you’re reviewing a scientific abstract. Your task is to identify the main entities (e.g., proteins, diseases) and their relations (e.g., causes, treats). For this task, this sentence : "input sentence" highlights the primary entity or relation in one word:"
@@ A.4 Prompts of the Sentiment Analysis Meta-Task
Sentiment Analysis Meta-Task Product Review Rating: In this task, you’re given a review from an online platform. Your task is to generate a rating for the product based on the review on a scale of 1-5, where 1 means ’extremely negative’ and 5 means ’extremely positive’. For this task, this sentence : "input sentence" reflects the sentiment in one word:" Emotion Detection: In this task, you’re reading a personal diary entry. Your task is to identify the predominant emotion expressed, such as joy, sadness, anger, fear, or love. For this task, this sentence : "input sentence" conveys the emotion in one word:" Sentiment Polarity Detection: In this task, you’re analyzing customer feedback from various platforms. Your task is to identify the overall sentiment polarity of the feedback. The sentiment polarity means: 1 for very negative, 2 for negative, 3 for neutral, 4 for positive, and 5 for very positive. Based on this guidance, this sentence : "input sentence" represents in one word:" Sentiment Intensity and Emotion Detection: In this task, your objective is to gauge the intensity and type of emotion conveyed in a piece of text, such as a social media post or a product review. This involves not just identifying whether the sentiment is positive or negative, but also understanding the strength of that sentiment and the specific emotions involved (e.g., joy, anger, sadness, surprise). For this task, this sentence : "input sentence" conveys an emotion that is best described in one word as:" Aspect-based Sentiment Analysis: In this task, you’re given a review of a product or service. Your task is to assess the sentiment toward specific aspects of the product or service mentioned in the review. For each mentioned aspect (e.g., quality, price, customer service), classify the sentiment as: 1 for very negative, 2 for negative, 3 for neutral, 4 for positive, and 5 for very positive. Based on this instruction, this sentence : "input sentence" signifies in one word:"
A.5 Task-Specific Prompts on Transfer Tasks
@@
MR/SST In this task, you’re given a movie review, and you need to classify its sentiment into positive or negative. For this task, this sentence : "input sentence" means in one word:" CR In this task, you’re given a customer review of a product sold online, and you need to classify its sentiment into positive or negative. For this task, this sentence : "input sentence" means in one word:" SUBJ In this task, you’re analyzing movie reviews to determine their level of subjectivity. A subjective review is filled with personal opinions, feelings, and preferences of the reviewer, often expressing likes or dislikes and personal experiences. An objective review, on the other hand, sticks to factual information, such as plot details or actor performances, without revealing the reviewer’s personal stance. For this task, this sentence : "input sentence" means in one word:" MPQA In this task, you are given a description of a entity or event expressed in data such as blogs, newswire, and editorials. You need to classify its sentiment into positive or negative. For this task, this sentence : "input sentence" means in one word:" TREC In this task, you are given a question. You need to detect which category better describes the question. A question belongs to the description category if it asks about description and abstract concepts. Entity questions are about entities such as animals, colors, sports, etc. Abbreviation questions ask about abbreviations and expressions abbreviated. Questions regarding human beings, description of a person, and a group or organization of persons are categorized as Human. Quantity questions are asking about numeric values and Location questions ask about locations, cities, and countries. Answer with "Description", "Entity", "Abbreviation", "Person", "Quantity", and "Location". For this task, this sentence : "input sentence" means in one word:" MRPC In this task, you are given two sentences(Sentence1 and Sentence2). Answer "Yes" if these sentences are a paraphrase of one another, otherwise answer "No". For this task, this sentence : "input sentence" means in one word:"

### v2

@@ 4.2 Main Results
The results of MetaEOL on STS tasks are shown in Table 1, with notable performance by MetaEOL which requires no training. Among models that do not require training, prompt-based methods exhibit superior results compared to average pooling methods, especially with the LLAMA and Mistral models. Across various models including LLAMA2-7B/13B, LLAMA3-8B, and Mistral-7B, MetaEOL demonstrates competitive performance compared to contrastive-trained models such as SimCSE-BERT and SimCSE-Roberta, albeit with a slight lag behind PromptBERT. Furthermore, MetaEOL significantly outperforms PromptEOL and Echo embeddings across various test models, demonstrating a consistent improvement. Notably, the LLAMA2-13B model using MetaEOL shows an average improvement of 6.73% over PromptEOL, underscoring the efficacy of MetaEOL. Compared to LLM2Vec which requires two-stage training, MetaEOL is competitive when using LLAMA2-7B, without the need for any training.
@@ 5.1 Ablation Study
We evaluate the effectiveness of key components of MetaEOL in Table 3. First, to ensure the improvement observed with MetaEOL is not merely due to involving more prompts, we create seven paraphrased versions of the PromptEOL prompt, resulting in a total of eight prompts. [^4] The seven paraphrased prompts are presented in Appendix A.2 We then average the embeddings from these eight prompts to form the final sentence embedding. We find merely duplicating PromptEOL prompts (w. 7 paraphrase) does not improve PromptEOL but results in a significant decline. Additionally, we implement MetaEOL exclusively on each meta-task (TC/SA/PI/IE only). We find that tasks requiring a detailed comprehension of sentences (PI and IE) yield superior performance compared to those requiring a broader understanding, even surpassing PromptEOL. MetaEOL, which combines the embeddings from these meta-tasks, outperforms all individual meta-tasks, confirming the complementarity of the meta-tasks and the effectiveness of combining embeddings from diverse meta-tasks. We finally find that averaging the embeddings from different meta-tasks yields better results than either concatenating them or max pooling them across each dimension.
@@ 5.3 Influence of Number of Prompts
5.4 Prompt Sensitivity Analysis
To test the sensitivity of MetaEOL, we specifically focus on the influence of (i) Tiny perturbations on the task prompt; and (ii) Variations of the major prompting instruction in Appendix A.1.
5.4.1 Sensitivity to Tiny Prompt Perturbations
We apply the synonym replacement operation in Wei and Zou (2019) to replace 10% of words in the sentiment analysis task prompt of MetaEOL with their synonyms. We craft an additional 4 perturbed prompts. The synonym replacements are sourced directly from WordNet without filtering, which often results in unnatural substitutions, as shown in the Appendix A.5. To provide context for the sensitivity, we included results from SimCSE-BERT-Base with varying random seeds in Jiang et al. (2022a) as a reference.
Method STS Avg.
MetaEOL-LLAMA2-7B 67.98±0.67
SimCSE-BERT-Base 75.42±0.86
Table 5: Sensitivity to tiny prompt perturbations on LLAMA2-7B.
The results in Table 5 show that even with unnatural substitutions, our MetaEOL-LLAMA2-7B still exhibits a standard deviation of ±0.67 on STS Avg., which is in line with the variance observed in SimCSE-BERT-Base (±0.86), suggesting that our method’s sensitivity to prompt perturbations is comparable to that of existing approaches to random seeds.
5.4.2 Sensitivity to Variations of the Major Prompting Instruction
We vary the example task prompt in the major prompting instruction in Appendix A.1 with task prompts from the four meta-tasks used in our MetaEOL. As the major prompting instruction is used to prompt ChatGPT-4 to generate task prompts, changing it will lead to a completely different set of task prompts.
Method STS Avg.
MetaEOL-LLAMA2-7B 76.17±1.06
SimCSE-BERT-Base 75.42±0.86
Table 6: Sensitivity to variations of the major prompting instruction on LLAMA2-7B.
Overall the results show MetaEOL with LLAMA2-7B can beat tuned SimCSE-BERT-Base on average, and is worse but still comparable to SimCSE-Bert-Base in terms of standard deviation, suggesting that MetaEOL’s sensitivity to major prompting instruction’s variations is also comparable to that of existing approaches to random seeds.
@@ 5.5 Influence of Output Layers
We check the impact of output layers for LLAMA2 and Mistral-7B models, using PromptEOL and MetaEOL. Figure 4 shows that the last layer is not always the most effective for STS tasks, which is consistent with the findings in Li and Li (2024b).
@@
5.6 Scaling LLMs
Previous study Jiang et al. (2023b) show scaling model sizes does not lead to performance improvement on STS tasks. In this section, we investigate the impact of model size on the performance of MetaEOL. For the sake of computational resources, we load models with 4-bit precision.
The results in Table 7 show that using the final layer for sentence embedding generation, which is indicated by layer index -1, does not yield improved performance with increased model size. Contrastingly, the application of our proportional layer strategy reveals a different trend. Specifically, the LLAMA2-70B model, which utilizes the -8 layer, demonstrates superior performance, suggesting that larger models might benefit more significantly from selecting a proportionate layer rather than the last layer for sentence embedding. This observation could point to a potential scaling law, where larger models require a different, non-final layer to maximize performance effectively.
5.7 Transfer Learning Tasks
We conclude our analysis by assessing the performance of MetaEOL on transfer learning tasks. Following prior works (Gao et al., 2021; Ni et al., 2022a), we utilize the standard transfer learning tasks provided by SentEval. The tasks consist of MR (Pang and Lee, 2005), CR (Hu and Liu, 2004), SUBJ (Pang and Lee, 2004), MPQA (Wiebe et al., 2005), SST-2 (Socher et al., 2013), TREC (Voorhees and Tice, 2000), and MRPC (Dolan and Brockett, 2005). For each task, logistic regression classifiers are trained using the created sentence embeddings as input features. The test accuracy on each task is used as the metric. Additionally, we include two supervised contrastive-trained models (SimCSE and ST5-Enc) for reference. Notably, ST5-Enc, a model with a 4.8B parameter count, is extensively trained on natural language inference (NLI) data and two billion question-answer pairs.
To investigate the ability of task-specific prompts to modify embedding behavior, we have crafted task prompts tailored to each SentEval task. [^6] The details of the task prompts are in Appendix A.6. As an example, for the Movie Review (MR) dataset, we designed a prompt structured as: In this task, you’re given a movie review, and you need to classify its sentiment into positive or negative. For this task, this sentence: "input sentence" means in one word:", referred to as MRPrompt in Table 8. These task-specific prompts significantly improve the corresponding task performance, always better than PromptEOL and heavily supervised contrastive-trained ST5-Enc, verifying that LLAMA2-7B can follow the prompt to generate tailored embeddings without any training. This indicates that carefully designed prompts can effectively steer the pre-trained embeddings to align with various NLP tasks, thus providing a more resource-efficient alternative to the traditional fine-tuning paradigm.
@@ 6 Conclusion
In this paper, we introduce MetaEOL, a new approach for deriving high-quality sentence embeddings from LLMs without requiring any training. By leveraging a diverse set of meta-task prompts, MetaEOL effectively captures multiple representations of sentences from distinct perspectives. We show simply averaging these meta-task derived embeddings leads to generalized general-purpose embeddings, which work remarkably well across STS datasets and transfer learning tasks.
@@
We note two limitations in our work: computational overhead and restricted evaluation benchmarks. As MetaEOL requires feeding multiple prompts to LLMs to generate several embeddings, the computational cost will be higher than that of previous methods. Our results indicate that increasing the number of tasks leads to performance improvements, but it also worsens the efficiency issue. If the number of prompts is increased, the efficiency of our approach would further decrease. Nonetheless, in contexts where sentences are consistently reused, such as when embeddings are stored for downstream classification or retrieval tasks, the issue becomes less significant. Our evaluation is currently confined to sentence-level tasks in English only. As LLMs continue to advance, exploring the performance of MetaEOL in multilingual contexts and its applicability to document retrieval Zhuang et al. (2024) presents an intriguing avenue for future research.
Acknowledgement
This research was supported by the Hybrid Intelligence Center, a 10-year program funded by the Dutch Ministry of Education, Culture and Science through the Netherlands Organisation for Scientific Research, https://hybrid-intelligence-centre.nl, and project VI.Vidi.223.166 of the NWO Talent Programme which is (partly) financed by the Dutch Research Council (NWO).
@@ A.2 Paraphrased Prompts of PromptEOL
1. This sentence : "input sentence" can be rephrased to one word:" 2. This sentence : "input sentence" can be expressed as one word:" 3. This sentence : "input sentence" implies in one word:" 4. This sentence : "input sentence" indicates in one word:" 5. The meaning of this sentence : "input sentence" can be conveyed in another word:" 6. This sentence : "input sentence" can be restated as one word:" 7. This sentence : "input sentence" can be reformulated as one word:"
@@ A.3 Prompts of MetaEOL
Text Classification
General Category Identification
: In this task, you’re presented with a text excerpt. Your task is to categorize the excerpt into a broad category such as ’Education’, ’Technology’, ’Health’, ’Business’, ’Environment’, ’Politics’, or ’Culture’. These categories help in organizing content for better accessibility and targeting. For this task, this sentence : "input sentence" should be classified under one general category in one word:"
Opinion vs. Fact Discrimination: In this task, you’re given a statement and you need to determine whether it’s presenting an ’Opinion’ or a ’Fact’. This distinction is vital for information verification, educational purposes, and content analysis. For this task, this sentence : "input sentence" discriminates between opinion and fact in one word:" Sentiment Analysis
Product Review Rating
: In this task, you’re given a review from an online platform. Your task is to generate a rating for the product based on the review on a scale of 1-5, where 1 means ’extremely negative’ and 5 means ’extremely positive’. For this task, this sentence : "input sentence" reflects the sentiment in one word:"
Emotion Detection: In this task, you’re reading a personal diary entry. Your task is to identify the predominant emotion expressed, such as joy, sadness, anger, fear, or love. For this task, this sentence : "input sentence" conveys the emotion in one word:" Paraphrase Identification
Similarity Check
: In this task, you’re presented with two sentences. Your task is to assess whether the sentences convey the same meaning. Use ’identical’, ’similar’, ’different’, or ’unrelated’ to describe the relationship. To enhance the performance of this task, this sentence : "input sentence" means in one word:"
Contextual Synonym Detection: In this task, you’re given a sentence and a phrase. Your task is to determine if the phrase can be a contextual synonym within the given sentence. Options include ’yes’, ’no’, or ’partially’. To enhance the performance of this task, this sentence : "input sentence" means in one word:" Information Extraction
Key Fact Identification
: In this task, you’re examining a news article. Your task is to extract the most critical fact from the article. For this task, this sentence : "input sentence" encapsulates the key fact in one word:"
Entity and Relation Extraction: In this task, you’re reviewing a scientific abstract. Your task is to identify the main entities (e.g., proteins, diseases) and their relations (e.g., causes, treats). For this task, this sentence : "input sentence" highlights the primary entity or relation in one word:"
@@ A.4 Prompts of the Sentiment Analysis Meta-Task
Sentiment Analysis Meta-Task
Product Review Rating
: In this task, you’re given a review from an online platform. Your task is to generate a rating for the product based on the review on a scale of 1-5, where 1 means ’extremely negative’ and 5 means ’extremely positive’. For this task, this sentence : "input sentence" reflects the sentiment in one word:"
@@
Emotion Detection: In this task, you’re reading a personal diary entry. Your task is to identify the predominant emotion expressed, such as joy, sadness, anger, fear, or love. For this task, this sentence : "input sentence" conveys the emotion in one word:" Sentiment Polarity Detection: In this task, you’re analyzing customer feedback from various platforms. Your task is to identify the overall sentiment polarity of the feedback. The sentiment polarity means: 1 for very negative, 2 for negative, 3 for neutral, 4 for positive, and 5 for very positive. Based on this guidance, this sentence : "input sentence" represents in one word:" Sentiment Intensity and Emotion Detection: In this task, your objective is to gauge the intensity and type of emotion conveyed in a piece of text, such as a social media post or a product review. This involves not just identifying whether the sentiment is positive or negative, but also understanding the strength of that sentiment and the specific emotions involved (e.g., joy, anger, sadness, surprise). For this task, this sentence : "input sentence" conveys an emotion that is best described in one word as:" Aspect-based Sentiment Analysis: In this task, you’re given a review of a product or service. Your task is to assess the sentiment toward specific aspects of the product or service mentioned in the review. For each mentioned aspect (e.g., quality, price, customer service), classify the sentiment as: 1 for very negative, 2 for negative, 3 for neutral, 4 for positive, and 5 for very positive. Based on this instruction, this sentence : "input sentence" signifies in one word:"
A.5 Sentiment Analysis Task Prompts with Tiny Perturbatios
Original
In this task, you’re given a review from an online platform. Your task is to generate a rating for the product based on the review on a scale of 1-5, where 1 means ’extremely negative’ and 5 means ’extremely positive’. For this task, this sentence : "input sentence" reflects the sentiment in one word:"
Perturbed
1.
In this task, you’re given a reappraisal from an online chopine. Your task is to generate a rating for the product based on the reappraisal on a scale of 1-5, where 1 think of ’extremely negative’ and 5 think of ’extremely positive’. For this task, this sentence : "input sentence" reflects the sentiment in one word:"
2. In this task, you’re given a review from an online chopine. Your task is to generate a rating for the product based on the review on a scale of 1-5, where 1 means ’extremely damaging’ and 5 means ’extremely plus’. For this task, this sentence : "input sentence" reflects the sentiment in one word:" 3. In this job, you’re given a brush up from an online platform. Your job is to generate a rating for the product based on the brush up on a scale of 1-5, where 1 means ’highly negative’ and 5 means ’highly positive’. For this task, this sentence : "input sentence" reflects the sentiment in one word:" 4. In this task, you’re reach a refresh from an online platform. Your task is to generate a rating for the product based on the refresh on a scale of 1-5, where 1 means ’highly negative’ and 5 means ’highly positive’. For this task, this sentence : "input sentence" reflects the sentiment in one word:"
A.6 Task-Specific Prompts on Transfer Tasks
MR/SST
In this task, you’re given a movie review, and you need to classify its sentiment into positive or negative. For this task, this sentence : "input sentence" means in one word:"
CR
In this task, you’re given a customer review of a product sold online, and you need to classify its sentiment into positive or negative. For this task, this sentence : "input sentence" means in one word:"
SUBJ
In this task, you’re analyzing movie reviews to determine their level of subjectivity. A subjective review is filled with personal opinions, feelings, and preferences of the reviewer, often expressing likes or dislikes and personal experiences. An objective review, on the other hand, sticks to factual information, such as plot details or actor performances, without revealing the reviewer’s personal stance. For this task, this sentence : "input sentence" means in one word:"
MPQA
In this task, you are given a description of a entity or event expressed in data such as blogs, newswire, and editorials. You need to classify its sentiment into positive or negative. For this task, this sentence : "input sentence" means in one word:"
TREC
@@ A.6 Task-Specific Prompts on Transfer Tasks
In this task, you are given a question. You need to detect which category better describes the question. A question belongs to the description category if it asks about description and abstract concepts. Entity questions are about entities such as animals, colors, sports, etc. Abbreviation questions ask about abbreviations and expressions abbreviated. Questions regarding human beings, description of a person, and a group or organization of persons are categorized as Human. Quantity questions are asking about numeric values and Location questions ask about locations, cities, and countries. Answer with "Description", "Entity", "Abbreviation", "Person", "Quantity", and "Location". For this task, this sentence : "input sentence" means in one word:"
MRPC
In this task, you are given two sentences(Sentence1 and Sentence2). Answer "Yes" if these sentences are a paraphrase of one another, otherwise answer "No". For this task, this sentence : "input sentence" means in one word:"